# MARS-M: When Variance Reduction Meets Matrices

### v2

@@ 3.2 MARS
A typical form for variance reduction is exemplified by STORM (Cutkosky and Orabona, 2019), which introduces gradient correction to the momentum update rule:
$\mathbf{m}_{t}=\beta\mathbf{m}_{t-1}+(1-\beta)\Big[\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\underbrace{\frac{\beta}{1-\beta}\big(\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big)}_{\text{gradient correction}}\Big].$
where $\beta>0$ is momentum parameter. Here the noise brought by the randomness of data can be canceled out by the stochastic gradient difference term $\frac{\beta}{1-\beta}(\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t}))$. Based on this term, the estimation of the true gradient for the parameter in the last step $\nabla F(\mathbf{x}_{t-1})$ can be transferred to the true gradient for the parameter in the current step $\nabla F(\mathbf{x}_{t})$.
Based on STORM, MARS (Yuan et al., 2025) (See Algorithm 1) introduces a scaling parameter to the gradient correction part such that the corrected gradient is calculated as:
@@
where $\gamma_{t}>0$ is the coefficient for scaling. Moreover, it also introduces gradient clipping on the corrected gradient for stabler training:
@@
For the ease of exposure, the MARS algorithm is summarized in Algorithm 1, where line 7 is the general formula for adaptive gradient methods such as AdamW (Loshchilov and Hutter, 2019), Lion (Chen et al., 2023) and Shampoo (Gupta et al., 2018), with approximated Hessian matrix $\mathbf{H}_{t}$. As will be seen later, our work can be seen as a tailored version of MARS for Muon/moonlight.
@@ 4.1 MARS-M
It is worth noting that MARS-M is similar to MARS-Shampoo in (Yuan et al., 2025), which is another instantiation of MARS to a simplified version of Shampoo. In detail, the update rule in MARS-Shampoo is instantiated as:
@@
In practice, SVD computation can be substituted by approximate numerical methods such as Newton iteration (Lakić, 1998; Higham, 2008; Anil et al., 2020; Bernstein and Newhouse, 2024) and Newton-Schulz iteration (Schulz, 1933; Higham, 2008). The main differences between MARS-M and MARS-Shampoo are that MARS-M scales $\mathbf{O}_{t}\approx\mathbf{U}_{t}\mathbf{V}_{t}^{\top}$ by $0.2\cdot\sqrt{\max(m,n)}$ and incorporates weight decay $\lambda\mathbf{X}_{t}$, both of which are essential for achieving superior performance according to Liu et al. (2025).
@@ 4.2 Approximated MARS-M
Since calculating the stochastic gradients twice requires much more computation, in practice, approximate version of MARS is usually utilized where $\nabla f(\mathbf{X}_{t-1},\bm{\xi}_{t})$ is replaced by $\nabla f(\mathbf{X}_{t-1},\bm{\xi}_{t-1})$. If clipping on $\mathbf{C}_{t}$ is ignored, lines 4 to 5 in Algorithm 2 can be approximated by:
$\mathbf{C}_{t}$ $=\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})+\gamma_{t}\bigg(\frac{\beta}{1-\beta}\bigg)\big(\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{X}_{t-1},\bm{\xi}_{t-1})\big),$~ (4.2)
$\mathbf{M}_{t}$ $=\beta\mathbf{M}_{t-1}+(1-\beta)\mathbf{C}_{t}.$
@@
By comparing (4.3), (4.4) with (3.5), (3.6), we can see that approximate MARS-M can be seen as a variant of Moonlight with adjusted momentum parameters $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(1-\gamma_{t})(1-\beta)/\beta$ and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{t}$ (in contrast to 1’s in moonlight).
@@
$\mathbf{C}_{t}$ $={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{1}{1-\beta}}\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1}\cdot\bigg(\frac{\beta}{1-\beta}\bigg)\big(\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{X}_{t-1},\bm{\xi}_{t-1})\big),$~ (4.5)
@@
Comparing (4.2) with (4.5), it can be seen that approximate version of MARS-M is Moonlight with stochastic gradient $\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})$ being scaled down by a factor of $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(1-\beta)$ and setting $\gamma_{t}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1}$. The experiments in Section 5 empirically demonstrate that this difference in momentum parameters leads to a noticeable impact on the loss.
@@ 4.3 Convergence Analysis
To analyze the convergence of MARS-M, similar to Yuan et al. (2025), we first make the following assumptions:
@@
Both assumptions are standard in the literature (Cutkosky and Orabona, 2019; Yuan et al., 2025). We have the following theorem, which guarantees the convergence of MARS-M in Algorithm 2.
@@
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla F(\mathbf{X}_{t})\|_{F}\leq\frac{2\sqrt{2LG}}{T^{1/3}}+\frac{2\sqrt{2LB\log(s+T)}}{T^{1/3}}+\frac{2G}{T^{1/3}}+\frac{2B}{T^{1/3}}\log(s+T)-\frac{\sqrt{2}}{4LT^{1/3}}\sum_{t=1}^{T}\frac{\widetilde{M}_{t+1}}{\sqrt{\eta_{t}}}.$
@@
We defer the proof of the above theorem to Appendix C. It is worth noting that time-varying momentum parameter $\beta_{t}$ is required for theoretical analysis, while it can be chosen as a constant in practice. Theorem 4.3 suggests that MARS-M can achieve a non-asymptotic convergence rate of $\mathcal{O}(T^{-1/3}\log(T))$, the same rate as general MARS as proved in Yuan et al. (2025). As a comparison, Li and Hong (2025)Shen et al. (2025) and Pethick et al. (2025) proved that Muon can achieve only $\mathcal{O}(T^{-1/4})$ convergence rate.
To sum up, Muon-M constitutes a new addition to the MARS family, broadening the framework to encompass more matrix-based optimizers with strong theoretical guarantees.
5 Experiments
5.1 LLM Experiments
We evaluate the performance of our algorithm with baseline algorithms [^2] For training efficiency, we use approximated MARS-M for LLM experiments, since the performances of exact version and approximate version are similar as discussed in Yuan et al. (2025). And we compare their differences in computer vision experiments., including Moonlight and AdamW, on language modeling tasks based on the nanoGPT (Karpathy, 2022) architecture and GPT-2 (Radford et al., 2019) model. We implement experiments with OpenWebText (Gokaslan et al., 2019) and FineWeb-Edu 100B (Lozhkov et al., 2024) dataset. For OpenWebText, the training and validation datasets contain around $9$ billion and $4.4$ million tokens, respectively; while the training and validation sets of FineWeb-Edu 100B are with 100 billion and 0.1 billion tokens. The experiments are conducted on scales of small (125M parameters), medium (355M parameters), large (770M parameters) as well as XL (1.5B parameters). Moreover, we disabled biases, and set the Dropout rate (Srivastava et al., 2014) to $0.0$. For experiments with small models, we utilize 16 NVIDIA H800 GPUs; and for models with larger sizes, we implement with 32 NVIDIA H800. Other hyper-parameters of training are listed in Appendix A.
5.2 Experiment Results
We show the zoomed-in curves for training and validation losses for large-size models trained on OpenWebText and XL-size models trained on FineWeb-Edu 100B datasets in Figures 3 and 1. Moreover, we also show the curves for XL-size models for all the training process in Figure 2. Other curves can be found in Appendix B. We also add the curves for the experiments trained with AdamW in Yuan et al. (2025), where the maximum learning rate is 6e46e-4 for XL-size model and 2e42e-4 for large-size model. It can be observed that the experiments trained with MARS-M display steady improvement on both training and validation losses over Moonlight. Additionally, although the loss for AdamW is lower in the middle of training due to a smaller maximum learning rate, model trained with MARS-M achieves lower losses than that of AdamW in the final phase. Since the learning rates for all the experiments trained with baseline optimizers results from grid search, it can be concluded that MARS-M can indeed improve the performance of large language model training.
Figure 1: The training and validation loss of XL-size models (1.5B) trained with different optimizers on the FineWeb-Edu 100B dataset.
Figure 2: The training and validation loss of XL-size models (1.5B) trained with different optimizers on the FineWeb-Edu 100B dataset.
Figure 3: The training and validation loss of large-size models (770M) trained with different optimizers on the OpenWebText dataset.
Furthermore, We also evaluate 0-shot and 2-shot performances of these optimizers on benchmarks including ARC (Yadav et al., 2019), HellaSwag (Zellers et al., 2019), OBQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), SciQ (Johannes Welbl, 2017), WinoGrande (Sakaguchi et al., 2020) and MMLU (Hendrycks et al., 2021), based on lm-evaluation-harness codebase (Gao et al., 2024). We only display the 2-shot performances for XL models in Table 1, and postpone other results in the Appendix B. It can be observed that MARS-M outperforms Moonlight on most of the benchmarks, showing that our algorithm can enhance the performance of pre-trained large language models on a wide range of downstream tasks.
@@ 5.3 Computer Vision Experiments
Besides language model task, we also conduct experiments on computer vision tasks with these optimizers on CIFAR-10 dataset (Krizhevsky et al., 2009) and ResNet-18 model (He et al., 2016). We set $\gamma=0.025$ for MARS-M and do grid search over $\{10^{-4},\cdots,10^{-2}\}$ for the learning rate. We set $\beta=0.95$ for all the optimizers. All the experiments are done with training batch size 128 on 1 NVIDIA A6000 GPU, with a total of 200 training epochs. Additionally, we use MultiStepLR scheduler so that the learning rate will decrease to 10% at 100th step and 1% at 150th step. The test loss and accuracy for MARS-M and MARS-M approximate, as well as Moonlight optimizer, are shown in Figure 4. It can be seen that MARS-M achieve better test accuracy and lower test loss over Moonlight and MARS-M-approx, validating the effect of variance reduction.
@@ 5.4 Ablation Study
[…]
We also implement ablation study on $\gamma$ in MARS-M to unearth the effect of scale of gradient correction on the performance. The experiments are done on small models trained with OpenWebText datasets, and the results are shown in Figure 5. It can be observed that the scale on gradient correction should be much smaller than 1, as usually taken in most of the variance reduction approaches (Cutkosky and Orabona, 2019; Nguyen et al., 2017a; Fang et al., 2018). However, the training and validation losses for experiments with $\gamma\in[0.005,0.025]$ are very close, showing that a small enough $\gamma$ is relatively robust to variation.
@@ 6 Conclusion
In this work, we have introduced MARS-M, a new algorithm that integrates MARS-based variance reduction with Muon/Moonlight to accelerate large language model training. Our algorithm integrates the advantages of variance reduction and matrix-based optimization. Through empirical experiments on language model and computer vision tasks, we demonstrate that models trained with MARS-M gain steady edges on loss and evaluation on downstream tasks over Moonlight baseline. In summary, our algorithm demonstrates an effective pathway for improving state-of-the-art matrix-based optimizers using variance reduction techniques.

### v3

@@ 3.2 MARS
A typical form of variance reduction is exemplified by STORM (STORM), which introduces a gradient-correction term in the momentum update rule:
$\mathbf{m}_{t}$ $=\beta\mathbf{m}_{t-1}+(1-\beta)\Big[\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$
$\qquad+\underbrace{\frac{\beta}{1-\beta}\big(\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big)}_{\text{gradient correction}}\Big].$
where $\beta>0$ is the momentum parameter. Here the noise brought by the randomness of data can be canceled out by the stochastic gradient difference term $\frac{\beta}{1-\beta}(\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t}))$. Based on this term, the estimation of the true gradient for the parameter in the last step $\nabla F(\mathbf{x}_{t-1})$ can be transferred to the true gradient for the parameter in the current step $\nabla F(\mathbf{x}_{t})$.
Based on STORM, MARS (yuan2024mars) (see Algorithm 1) introduces a scaling parameter for the gradient-correction term, yielding the corrected gradient:
@@
where $\gamma_{t}>0$ is the scaling coefficient. Moreover, it also applies gradient clipping to the corrected gradient to improve training stability:
@@
For the ease of exposure, the MARS algorithm is summarized in Algorithm 1, where line 7 is the general formula for adaptive gradient methods such as AdamW (Adamw), Lion (Lion) and Shampoo (gupta2018shampoo), with an approximated Hessian matrix $\mathbf{H}_{t}$. As will be seen later, our work can be viewed as a tailored version of MARS for Muon/Moonlight.
@@ 4.1 MARS-M
It is worth noting that MARS-M is related to MARS-Shampoo in (yuan2024mars), which is another instantiation of MARS based on a simplified version of Shampoo. Specifically, the update rule in MARS-Shampoo is instantiated as:
@@
In practice, the SVD computation can be substituted by approximate numerical methods such as Newton iteration (lakic1998computation; higham2008functions; anil2020scalable; bernstein2024old) and Newton-Schulz iteration (schulz1933iterative; higham2008functions). The main differences between MARS-M and MARS-Shampoo are that MARS-M scales $\mathbf{O}_{t}\approx\mathbf{U}_{t}\mathbf{V}_{t}^{\top}$ by $0.2\cdot\sqrt{\max(m,n)}$ and incorporates weight decay $\lambda\mathbf{X}_{t}$, both of which are essential for achieving superior performance according to liu2025muon.
@@ 4.2 Approximated MARS-M
Since computing the stochastic gradients twice is significantly more expensive, in practice, an approximate version of MARS is usually utilized where $\nabla f(\mathbf{X}_{t-1},\bm{\xi}_{t})$ is replaced by $\nabla f(\mathbf{X}_{t-1},\bm{\xi}_{t-1})$. If clipping on $\mathbf{C}_{t}$ is ignored, lines 4 to 5 in Algorithm 2 can be approximated by:
$\mathbf{C}_{t}$ $=\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})$
$\qquad+\gamma_{t}\bigg(\frac{\beta}{1-\beta}\bigg)\big(\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{X}_{t-1},\bm{\xi}_{t-1})\big),$~ (4.2)
$\mathbf{M}_{t}$ $=\beta\mathbf{M}_{t-1}+(1-\beta)\mathbf{C}_{t}.$~ (4.3)
@@
By comparing (4.4) and (4.5) with (3.5) and (3.6), we can see that approximate MARS-M can be viewed as a variant of Moonlight with adjusted momentum parameters $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(1-\gamma_{t})(1-\beta)/\beta$ and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{t}$ (in contrast to the coefficients equal to 1 in Moonlight).
@@
$\mathbf{C}_{t}$ $={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{1}{1-\beta}}\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})$
$\qquad+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1}\cdot\bigg(\frac{\beta}{1-\beta}\bigg)\big(\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{X}_{t-1},\bm{\xi}_{t-1})\big),$~ (4.6)
@@
Comparing (4.2) with (4.6), it can be seen that the approximate version of MARS-M corresponds to Moonlight with the stochastic gradient $\nabla f(\mathbf{X}_{t},\bm{\xi}_{t})$ being scaled down by a factor of $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(1-\beta)$ and setting $\gamma_{t}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1}$. The experiments in Section 5 empirically demonstrate that this difference in momentum parameters leads to a noticeable impact on the loss.
@@ 4.3 Convergence Analysis
To analyze the convergence of MARS-M, similar to yuan2024mars, we first make the following assumptions:
@@
Both assumptions are standard in the literature (STORM; yuan2024mars). We have the following theorem, which guarantees the convergence of MARS-M in Algorithm 2.
@@
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla F(\mathbf{X}_{t})\|_{F}\leq$ $\frac{2\sqrt{2LG}}{T^{1/3}}+\frac{2\sqrt{2LB\log(s+T)}}{T^{1/3}}$
$\qquad+\frac{2G}{T^{1/3}}+\frac{2B}{T^{1/3}}\log(s+T)$
$\qquad-\frac{\sqrt{2}}{4LT^{1/3}}\sum_{t=1}^{T}\frac{\widetilde{M}_{t+1}}{\sqrt{\eta_{t}}}.$
@@
We defer the proof of the above theorem to Appendix C. It is worth noting that time-varying momentum parameter $\beta_{t}$ is required for theoretical analysis, while it can be chosen as a constant in practice. Theorem 4.3 suggests that MARS-M can achieve a non-asymptotic convergence rate of $\mathcal{O}(T^{-1/3}\log(T))$, the same rate as general MARS as proved in yuan2024mars. As a comparison, li2025noteshen2025convergence and pethick2025training proved that Muon can achieve only $\mathcal{O}(T^{-1/4})$ convergence rate.
To sum up, MARS-M constitutes a new addition to the MARS family, broadening the framework to encompass more matrix-based optimizers with strong theoretical guarantees.
Method ARC-C ARC-E Hellaswag MMLU OpenBookQA PIQA SciQ WG Avg.
AdamW 33.53 66.46 45.02 25.32 35.40 69.10 86.70 55.80 52.17
Method ARC-C ARC-E Hellaswag MMLU OpenBookQA PIQA SciQ WG Avg.
AdamW 38.05 70.29 50.30 26.87 38.20 70.46 92.10 55.80 55.26
@@
5 Experiments
5.1 LLM Experiments
We evaluate the performances of our algorithm with baseline algorithms [^2] For training efficiency, we use approximated MARS-M for the LLM experiments, since the performances of exact version and approximate versions performs similarly as discussed in yuan2024mars. And we compare their differences in computer vision experiments., including Moonlight and AdamW, on language modeling tasks based on the nanoGPT (Karpathy2022) architecture and GPT-2 (radford2019language) model. We conduct experiments on OpenWebText (Gokaslan2019OpenWeb) and FineWeb-Edu 100B (lozhkov2024fineweb-edu) datasets. For OpenWebText, the training and validation datasets contain approximately $9$ billion and $4.4$ million tokens, respectively; while the training and validation sets of FineWeb-Edu 100B are with 100 billion and 0.1 billion tokens. We train for 100,000 steps with 2,000 warm-up steps, using a context length of 1024 and a total batch size of 480.
We run experiments at four scales: small (125M parameters), medium (355M parameters), large (770M parameters) as well as XL (1.5B parameters). Additionally, we disable biases and set the dropout rate (srivastava2014dropout) to $0.0$. And we also use Cosine learning rate scheduler and set the gradient clipping threshold to 1.0. For training parameters, for experiments with either OpenWebText or FineWeb-Edu 100B datasets, we perform a grid search over learning rates between $\{5e-4,1e-3,3e-3,5e-3,6e-3,1e-2\}$ for Moonlight optimizer (the detailed learning rates are listed in Table 5 in Appendix A). We just apply the same learning rate for MARS-M optimizer. And we use $\beta=0.95$ for both Moonlight and MARS-M optimizers. Since Moonlight and MARS-M optimizers are designed only for matrix parameters, we optimize the vector-like parameters and embeddings with AdamW with the same learning rate.
For experiments with small models, we use 16 NVIDIA H800 GPUs; and for large models, we implement with 32 NVIDIA H800. Other training hyper-parameters are listed in Appendix A.
5.2 Experiment Results
We show the zoomed-in curves of training and validation losses for different sizes of models on OpenWebText and FineWeb-Edu 100B datasets in Figures 1-2 and Figures 3-4, respectively. And the entire curves can be found in Appendix B. We also plot the curves for the experiments trained with AdamW in yuan2024mars, and the corresponding learning rates is listed in Table 5, which are also attained from grid research as described in yuan2024mars.
It can be observed that the experiments trained with MARS-M display steady improvement on both training and validation losses over Moonlight. Additionally, although the loss for AdamW is lower in the middle of training due to a smaller maximum learning rate, model trained with MARS-M achieves lower losses than that of AdamW in the final phase. Since the learning rates for all the experiments trained with baseline optimizers results from grid search, it can be concluded that MARS-M can indeed improve the performance of large language model training.
Figure 1: The zoomed-in training loss of small-size (125M), medium-size (355M) and large-size (770M) models trained with different optimizers on the OpenWebText dataset.
Figure 2: The zoomed-in training and validation loss of large-size models (770M) trained with different optimizers on the OpenWebText dataset.
Figure 3: The zoomed-in training loss of medium-size (355M), large-size (770M) and XL-size (1.5B) models trained with different optimizers on the FineWeb-Edu 100B dataset.
Figure 4: The zoomed-in validation loss of medium-size (355M), large-size (770M) and XL-size (1.5B) models trained with different optimizers on the FineWeb-Edu 100B dataset.
Furthermore, We also evaluate 0-shot and 2-shot performances of these optimizers on benchmarks including ARC (ARC), HellaSwag (hellaswag), OBQA (OpenBookQA), PIQA (PIQA), SciQ (SciQ), WinoGrande (WinoGrande) and MMLU (MMLU), based on lm-evaluation-harness codebase (eval-harness). We only display the 2-shot performances for medium, large and XL models trained on FineWeb-Edu 100B dataset in Tables 12 and 3, respectively. And other results are postponed to the Appendix B. It can be observed that MARS-M outperforms Moonlight on most of the benchmarks, showing that our algorithm can enhance the performance of pre-trained large language models on a wide range of downstream tasks.
[…]
@@ 5.3 Computer Vision Experiments
Besides language model task, we also conduct experiments on computer vision tasks with these optimizers on CIFAR-10 dataset (krizhevsky2009learning) and ResNet-18 model (he2016deep). We set $\gamma=0.025$ for MARS-M and do grid search over $\{10^{-4},\cdots,10^{-2}\}$ for the learning rate. We set $\beta=0.95$ for all the optimizers. All the experiments are done with a training batch size of 128 on 1 NVIDIA A6000 GPU, with a total of 200 training epochs. Additionally, we use MultiStepLR scheduler so that the learning rate will decrease to 10% at 100th epoch and 1% at 150th epoch.
The test loss and accuracy for MARS-M and MARS-M approximate, as well as Moonlight optimizer, are shown in Figure 5. It can be seen that MARS-M achieve better test accuracy and lower test loss over Moonlight and MARS-M-approx, validating the effect of variance reduction.
@@ 5.4 Ablation Study
We also implement an ablation study on $\gamma$ in MARS-M to investigate how the scale of gradient-correction term affects performance. The experiments are performed on small models trained on OpenWebText datasets, and the results are shown in Figure 6. It can be observed that the scale on gradient correction should be much smaller than 1, as usually taken in most of the variance reduction approaches (STORM; SARAH1; fang2018spider). However, the training and validation losses for experiments with $\gamma\in[0.005,0.025]$ are very close, suggesting that a small enough $\gamma$ is relatively robust to variation.
@@ 6 Conclusion
In this work, we have introduced MARS-M, a new algorithm that integrates MARS-based variance reduction with Muon/Moonlight to accelerate large language model training. Our algorithm integrates the benefits of variance reduction and matrix-based optimization. Through empirical experiments on language model and computer vision tasks, we demonstrate that models trained with MARS-M gain steady edges on loss and evaluation on downstream tasks over Moonlight baseline. In summary, our algorithm demonstrates an effective pathway for improving state-of-the-art matrix-based optimizers using variance reduction techniques.